OpenStudy (anonymous):

if -q1q2/2 = (q1+q2)^2 then find q1/q2;

OpenStudy (anonymous):

\[-\frac{q_1 q_2}2 = q_1^2 + q_2^2 + 2q_1 q_2\] \[-\frac12 = \frac {q_1}q_2 + \frac{1}{\frac{q_1}{q_2}} + 2\] Let \(\frac{q_1}q_2 = k\) and solve it.

OpenStudy (anonymous):

You will get a quadratic equations in terms of k.

OpenStudy (anonymous):

hey yeah i got it as -3/4 :

OpenStudy (anonymous):

sorry no thats not what i got :

OpenStudy (anonymous):

i got -2 or -1/2

OpenStudy (anonymous):

0=k^2 + 1 + 5/2k, assuming k isn't zero 2k^2 + 2 + 5k = 0 http://www.wolframalpha.com/input/?i=2k%5E2+%2B+2+%2B+5k+%3D+0+ -2 and -1/2. so you're right.

OpenStudy (anonymous):

thanks a lot.